# Large Language Models for Information Retrieval: A Survey

### v2

@@ Abstract
As a primary means of information acquisition, information retrieval (IR) systems, such as search engines, have integrated themselves into our daily lives. These systems also serve as components of dialogue, question-answering, and recommender systems. The trajectory of IR has evolved dynamically from its origins in term-based methods to its integration with advanced neural models. While the neural models excel at capturing complex contextual signals and semantic nuances, thereby reshaping the IR landscape, they still face challenges such as data scarcity, interpretability, and the generation of contextually plausible yet potentially inaccurate responses. This evolution requires a combination of both traditional methods (such as term-based sparse retrieval methods with rapid response) and modern neural architectures (such as language models with powerful language understanding capacity). Meanwhile, the emergence of large language models (LLMs), typified by ChatGPT and GPT-4, has revolutionized natural language processing due to their remarkable language understanding, generation, generalization, and reasoning abilities. Consequently, recent research has sought to leverage LLMs to improve IR systems. Given the rapid evolution of this research trajectory, it is necessary to consolidate existing methodologies and provide nuanced insights through a comprehensive overview. In this survey, we delve into the confluence of LLMs and IR systems, including crucial aspects such as query rewriters, retrievers, rerankers, and readers. Additionally, we explore promising directions within this expanding field.
@@ 1 Introduction
Figure 1: Overview of existing studies that apply LLMs into information retrieval. LLMs can be used in query rewriter, retriever, reranker, and reader.
@@
The core function of an IR system is retrieval, which aims to determine the relevance between user-issued queries and the content to be retrieved, including various types of information such as texts, images, music, and more. For the scope of this survey, we concentrate solely on reviewing those text retrieval systems, in which query-document relevance is commonly measured by their matching score. [^3] The term “document” will henceforth refer to any text-based content subject to retrieve, including both long articles and short passages. Given that IR systems operate on extensive repositories, the efficiency of retrieval algorithms becomes of paramount importance. To improve the user experience, the retrieval performance is enhanced from both the upstream (query reformulation) and downstream (reranking and reading) perspectives. As an upstream technique, query reformulation is designed to refine user queries so that they are more effective at retrieving relevant documents [10, 11]. With the recent surge in the popularity of conversational search, this technique has received increasing attention. On the downstream side, reranking approaches are developed to further adjust the document ranking [12, 13, 14]. In contrast to the retrieval stage, reranking is performed only on a limited set of relevant documents, already retrieved by the retriever. Under this circumstance, the emphasis is placed on achieving higher performance rather than keeping higher efficiency, allowing for the application of more complex approaches in the reranking process. Additionally, reranking can accommodate other specific requirements, such as personalization [15, 16, 17, 18] and diversification [19, 20, 21, 22]. Following the retrieval and reranking stages, a reading component is incorporated to summarize the retrieved documents and deliver a concise document to users [23, 24]. While traditional IR systems typically require users to gather and organize relevant information themselves; however, the reading component is an integral part of new IR systems such as New Bing, [^4] New Bing, https://www.bing.com/new streamlining users’ browsing experience and saving valuable time.
@@
The trajectory of information retrieval (IR) has traversed a dynamic evolution, transitioning from its origins in term-based methods to the integration of neural models. Initially, IR was anchored in term-based methods [25] and Boolean logic, focusing on keyword matching for document retrieval. The paradigm gradually shifted with the introduction of vector space models [26], unlocking the potential to capture nuanced semantic relationships between terms. This progression continued with statistical language models [27, 28], refining relevance estimation through contextual and probabilistic considerations. The influential BM25 algorithm [29] played an important role during this phase, revolutionizing relevance ranking by accounting for term frequency and document length variations. The most recent chapter in IR’s journey is marked by the ascendancy of neural models [30, 3, 31, 32]. These models excel at capturing intricate contextual cues and semantic nuances, reshaping the landscape of IR. However, these neural models still face challenges such as data scarcity, interpretability, and the potential generation of plausible yet inaccurate responses. Thus, the evolution of IR continues to be a journey of balancing traditional strengths (such as the BM25 algorithm’s high efficiency) with the remarkable capability (such as semantic understanding) brought about by modern neural architectures.
@@
Our survey provides an insightful exploration of the intersection between LLMs and IR systems, covering key perspectives such as query rewriters, retrievers, rerankers, and readers (as shown in Figure 1). [^5] As yet, there has not been a formal definition for LLMs. In this paper, we mainly focus on models with more than 1B parameters. We also notice that some methods do not rely on such strictly defined LLMs, but due to their representativeness, we still include an introduction to them in this survey. This analysis enhances our understanding of LLMs’ potential and limitations in advancing the IR field. For this survey, we create a Github repository by collecting the relevant papers and resources about LLM4IR. [^6] https://github.com/RUC-NLPIR/LLM4IR-Survey We will continue to update the repository with newer papers. This survey will also be periodically updated according to the development of this area. We notice that there are several surveys for PLMs, LLMs, and their applications (e.g., AIGC or recommender systems) [46, 47, 48, 49, 50, 51, 52]. Among these, we highly recommend the survey of LLMs [52], which provides a systematic and comprehensive reference to many important aspects of LLMs. Compared with them, we focus on the techniques and methods for developing and applying LLMs for IR systems. In addition, we notice a perspective paper discussing the opportunity of IR when meeting LLMs [53]. It would be an excellent supplement to this survey regarding future directions.
The remaining part of this survey is organized as follows: Section 2 introduces the background for IR and LLMs. Section 345V respectively review recent progress from the four perspectives of query rewriter, retriever, reranker, and reader, which are four key components of an IR system. Then, Section 7 discusses some potential directions in future research. Finally, we conclude the survey in Section 8 by summarizing the major findings.
@@ 2.1 Information Retrieval
Researchers have identified several challenges with implications for the performance and effectiveness of IR systems, such as query ambiguity and retrieval efficiency. In light of these challenges, researchers have directed their attention towards crucial modules within the retrieval process, aiming to address specific issues and effectuate corresponding enhancements. The pivotal role of these modules in ameliorating the IR pipeline and elevating system performance cannot be overstated. In this survey, we focus on the following four modules, which have been greatly enhanced by LLMs.
Query Rewriter is an essential IR module that seeks to improve both the precision and expressiveness of user queries. Positioned at the early stage of the IR pipeline, this module assumes the crucial role of refining or modifying the initial query to align more accurately with the user’s information requirements. As an integral part of query rewriting, query expansion techniques, with pseudo relevance feedback being a prominent example, represent the mainstream approach to achieving query expression refinement. In addition to its utility in improving search effectiveness across general scenarios, the query rewriter finds application in diverse specialized retrieval contexts, such as personalized search and conversational search, thus further demonstrates its significance.
@@
Reader has evolved as a crucial module with the rapid development of large language model technologies. Its ability to comprehend real-time user intent and generate dynamic responses based on the retrieved text has revolutionized the presentation of IR results. In comparison to presenting a list of candidate documents, the reader module organizes answer texts in a more intuitive manner, simulating the natural way humans access information. To enhance the credibility of generated responses, the integration of references into generated responses has been an effective technique of the reader module.
@@ 2.2 Large Language Models
The evolution of LMs can be categorized into four primary stages, as discussed in prior literature [52]. Initially, LMs were rooted in statistical learning techniques and were termed statistical language models. These models tackled the issue of word prediction by employing the Markov assumption to predict the subsequent word based on preceding words. Thereafter, neural networks, particularly recurrent neural networks (RNNs), were introduced to calculate the likelihood of text sequences and establish neural language models. These advancements made it feasible to utilize LMs for representation learning beyond mere word sequence modeling. ELMo [58] first proposed to learn contextualized word representations through pre-training a bidirectional LSTM (biLSTM) network on large-scale corpora, followed by fine-tuning on specific downstream tasks. Similarly, BERT [59] proposed to pre-train a Transformer [60] encoder with a specially designed Masked Language Modeling (MLM) task and Next Sentence Prediction (NSP) task on large corpora. These studies initiated a new era of pre-trained language models (PLMs), with the “pre-training then fine-tuning” paradigm emerging as the prevailing learning approach. Along this line, numerous generative PLMs (e.g., GPT-2 [33], BART [61], and T5 [57]) have been developed for text generation problems including summarization, machine translation, and dialogue generation. Recently, researchers have observed that increasing the scale of PLMs (e.g., model size or data amount) can consistently improve their performance on downstream tasks (a phenomenon commonly referred to as the scaling law [62, 63]). Moreover, large-sized PLMs exhibit promising abilities (termed emergent abilities [42]) in addressing complex tasks, which are not evident in their smaller counterparts. Therefore, the research community refers to these large-sized PLMs large language models (LLMs).
@@ 3 Query Rewriter
In traditional IR systems, the user inputs a query, and the system returns a list of documents matching the query terms. However, original queries are often short or ambiguous, making them more susceptible to the problem of vocabulary mismatch. For example, if a user inputs the query “automobile” into a search engine, they expect to find information about cars. Nevertheless, the majority of relevant documents in the search results use the term “vehicle” rather than “automobile”. Since the search engine primarily relies on the exact query terms, it may not retrieve the most relevant documents containing the term “vehicle”, leading to a vocabulary mismatch challenge. Furthermore, in modern forms of IR systems, such as conversational search, the task of query rewriting plays an even more crucial role.
Traditional query rewriting methods have been developed to tackle this challenge by iteratively refining the user’s original query based on an analysis of the top-ranked documents retrieved by the initial query. These methods have demonstrated their effectiveness in improving retrieval accuracy by iteratively refining the query representation and incorporating additional information from the document collection. Notable examples of such methods include RM3 [87], LCE [88], KL expansion [89], and relevance modeling [90]. However, these methods often rely on predefined rules or heuristics, limiting their ability to fully capture the subtle nuances of user intent.
Fortunately, recent advancements in LLMs present promising opportunities to augment query rewriting capabilities and overcome the limitations of traditional approaches. These powerful language models have the potential to leverage vast amounts of textual data and learn more nuanced representations of queries and documents. By leveraging the capabilities of these models, researchers can explore novel approaches to query rewriting that can better align with the complex and diverse information needs of users. In the following sections, we will introduce the details of employing LLMs in query rewriting.
Figure 4: An example of LLM-based query rewriting for conversational search. The example is cited from LLMCS [91]. The LLM is used to generate a query based on the demonstrations and previous search context. Additional responses are required to generate for improving the query understanding. $N=0$ and $N>0$ correspond to zero-shot and few-shot scenarios.
3.1 Re-writing Target
Query rewriting typically serves two scenarios: ad-hoc retrieval, which addresses vocabulary mismatches between queries and documents, and conversational search, which refines and adapts system responses based on evolving conversations. In the following section, we will elaborate on the motivations behind these pursuits.
@@ 3.1.1 Ad-hoc Retrieval
Query rewriting plays a crucial role in ad-hoc retrieval as it serves as the initial step in the search funnel, and the overall effectiveness of the search process highly relies on the quality of query rewriting results. Within the context of ad-hoc retrieval, the primary objective of query rewriting is to retrieve a collection of documents that align more closely with the user’s information needs. The utilization of large language models (LLMs) in query rewriting offers several notable advantages, which can be summarized as follows:
$\bullet$ Better Semantic Understanding. LLMs have a deep understanding of language semantics, allowing them to capture the meaning and context of queries more effectively. As a comparison, traditional query rewriting methods primarily rely on statistical analysis of term frequencies [87, 88, 89], which may not capture the real intent of the query.
$\bullet$ Broad Knowledge. LLMs possess extensive knowledge, allowing them to draw from a wide range of concepts, facts, and information. This knowledge enables them to generate relevant and contextually appropriate query rewrites by leveraging their understanding of various topics.
$\bullet$ Independence of First-pass Retrieval. Traditional pseudo relevance feedback (PRF) methods retrieve a set of pseudo relevant documents as a source to refine the original query. However, the presence of irrelevant results in the pseudo relevance feedback set can introduce noise and potentially damage the retrieval performance. In contrast, LLMs could generate query rewrites directly based on the original query, which is independent of the first-pass retrieval and prevents the potential noise.
Query2Doc [86] is an LLM-base query rewriter, which prompts the LLMs to generate a relevant passage (which is equal to generate answers to some extent) according to the query. Subsequently, the original query is expanded by incorporating the generated passage. As shown in Figure 3, the original query written by the user is “when was Pokemon Green released”. The LLM generates a relevant passage of the original query by employing a few-shot prompting paradigm. Then, the original query and the generated passage are combined to construct a new query. The retriever module uses this new query to retrieve a list of relevant documents. Notably, the generated passage contains additional detailed information, such as “Pokemon Green was released in Japan on February 27th”, which effectively mitigates the vocabulary mismatch issue to some extent.
@@ 3.1.2 Conversational Search
Conversational search, on the other hand, involves a dynamic interaction between users and the retrieval system, where the system responds to user queries and clarifies the user’s information needs through a series of back-and-forth exchanges. In this context, query rewriting serves the purpose of refining and adapting the system’s responses based on the evolving conversation. The system may need to rewrite the user’s queries or generate new queries to retrieve relevant information that aligns with the ongoing conversation.
@@
To leverage the contextual understanding and generation capabilities of LLMs in conversational search, a framework called LLMCS (“Large Language Models know your Contextual Search intent”) has been proposed [91]. As shown in Figure 4, LLMCS prompts LLMs to generate query rewrites and longer hypothetical system responses from multiple perspectives. These generated outputs are then aggregated into an integrated representation that robustly captures the user’s complete search intent. Experimental results demonstrate that incorporating hypothetical responses alongside concise query rewrites significantly enhances search performance by explicitly supplementing more plausible search intent.
3.2 Data Resources
Query rewriting methods commonly require supplementary corpora to enrich the original queries. LLMs inherently incorporate world knowledge through their parameters. However, such general knowledge might prove inadequate for specific domains, thus necessitating domain-specific corpora to provide more comprehensive and domain-specific information. In this section, we will analyze two approaches: LLM-only methods that rely solely on the pre-existing knowledge within the model and corpus-enhanced LLM-based methods that leverage domain-specific corpora to augment the LLM’s capabilities.
3.2.1 Inherent Knowledge of LLMs

### v3

@@ Abstract
As a primary means of information acquisition, information retrieval (IR) systems, such as search engines, have integrated themselves into our daily lives. These systems also serve as components of dialogue, question-answering, and recommender systems. The trajectory of IR has evolved dynamically from its origins in term-based methods to its integration with advanced neural models. While the neural models excel at capturing complex contextual signals and semantic nuances, thereby reshaping the IR landscape, they still face challenges such as data scarcity, interpretability, and the generation of contextually plausible yet potentially inaccurate responses. This evolution requires a combination of both traditional methods (such as term-based sparse retrieval methods with rapid response) and modern neural architectures (such as language models with powerful language understanding capacity). Meanwhile, the emergence of large language models (LLMs), typified by ChatGPT and GPT-4, has revolutionized natural language processing due to their remarkable language understanding, generation, generalization, and reasoning abilities. Consequently, recent research has sought to leverage LLMs to improve IR systems. Given the rapid evolution of this research trajectory, it is necessary to consolidate existing methodologies and provide nuanced insights through a comprehensive overview. In this survey, we delve into the confluence of LLMs and IR systems, including crucial aspects such as query rewriters, retrievers, rerankers, and readers. Additionally, we explore promising directions, such as search agents, within this expanding field.
@@ 1 Introduction
Figure 1: Overview of existing studies that apply LLMs into IR. (1) LLMs can be used to enhance traditional IR components, such as query rewriter, retriever, reranker, and reader. (2) LLMs can also be used as search agents to perform multiple IR tasks.
@@
The core function of an IR system is retrieval, which aims to determine the relevance between a user-issued query and the content to be retrieved, including various types of information such as texts, images, music, and more. For the scope of this survey, we concentrate solely on reviewing those text retrieval systems, in which query-document relevance is commonly measured by their matching score. [^3] The term “document” will henceforth refer to any text-based content subject to retrieve, including both long articles and short passages. Given that IR systems operate on extensive repositories, the efficiency of retrieval algorithms becomes of paramount importance. To improve the user experience, the retrieval performance is enhanced from both the upstream (query reformulation) and downstream (reranking and reading) perspectives. As an upstream technique, query reformulation is designed to refine user queries so that they are more effective at retrieving relevant documents [10, 11]. With the recent surge in the popularity of conversational search, this technique has received increasing attention. On the downstream side, reranking approaches are developed to further adjust the document ranking [12, 13, 14]. In contrast to the retrieval stage, reranking is performed only on a limited set of relevant documents, already retrieved by the retriever. Under this circumstance, the emphasis is placed on achieving higher performance rather than keeping higher efficiency, allowing for the application of more complex approaches in the reranking process. Additionally, reranking can accommodate other specific requirements, such as personalization [15, 16, 17, 18] and diversification [19, 20, 21, 22]. Following the retrieval and reranking stages, a reading component is incorporated to summarize the retrieved documents and deliver a concise document to users [23, 24]. While traditional IR systems typically require users to gather and organize relevant information themselves; however, the reading component is an integral part of new IR systems such as New Bing, [^4] New Bing, https://www.bing.com/new streamlining users’ browsing experience and saving valuable time.
@@
The trajectory of IR has traversed a dynamic evolution, transitioning from its origins in term-based methods to the integration of neural models. Initially, IR was anchored in term-based methods [25] and Boolean logic, focusing on keyword matching for document retrieval. The paradigm gradually shifted with the introduction of vector space models [26], unlocking the potential to capture nuanced semantic relationships between terms. This progression continued with statistical language models [27, 28], refining relevance estimation through contextual and probabilistic considerations. The influential BM25 algorithm [29] played an important role during this phase, revolutionizing relevance ranking by accounting for term frequency and document length variations. The most recent chapter in IR’s journey is marked by the ascendancy of neural models [30, 3, 31, 32]. These models excel at capturing intricate contextual cues and semantic nuances, reshaping the landscape of IR. However, these neural models still face challenges such as data scarcity, interpretability, and the potential generation of plausible yet inaccurate responses. Thus, the evolution of IR continues to be a journey of balancing traditional strengths (such as the BM25 algorithm’s high efficiency) with the remarkable capability (such as semantic understanding) brought about by modern neural architectures.
@@
Our survey provides an insightful exploration of the intersection between LLMs and IR systems, covering key perspectives such as query rewriters, retrievers, rerankers, and readers (as shown in Figure 1). [^5] As yet, there has not been a formal definition for LLMs. In this paper, we mainly focus on models with more than 1B parameters. We also notice that some methods do not rely on such strictly defined LLMs, but due to their representativeness, we still include an introduction to them in this survey. We also include some recent studies that leverage LLMs as search agents to perform various IR tasks. This analysis enhances our understanding of LLMs’ potential and limitations in advancing the IR field. For this survey, we create a Github repository by collecting the relevant papers and resources about LLM4IR. [^6] https://github.com/RUC-NLPIR/LLM4IR-Survey We will continue to update the repository with newer papers. This survey will also be periodically updated according to the development of this area. We notice that there are several surveys for PLMs, LLMs, and their applications (e.g., AIGC or recommender systems) [46, 47, 48, 49, 50, 51, 52]. Among these, we highly recommend the survey of LLMs [52], which provides a systematic and comprehensive reference to many important aspects of LLMs. Compared with them, we focus on the techniques and methods for developing and applying LLMs for IR systems. In addition, we notice a perspective paper discussing the opportunity of IR when meeting LLMs [53]. It would be an excellent supplement to this survey regarding future directions.
The remaining part of this survey is organized as follows: Section 2 introduces the background for IR and LLMs. Section 345VII respectively review recent progress from the four perspectives of query rewriter, retriever, reranker, and reader, which are four key components of an IR system. Then, Section 8 discusses some potential directions in future research. Finally, we conclude the survey in Section 9 by summarizing the major findings.
@@ 2.1 Information Retrieval
Researchers have identified several challenges with implications for the performance and effectiveness of IR systems, such as query ambiguity and retrieval efficiency. In light of these challenges, researchers have directed their attention toward crucial modules within the retrieval process, aiming to address specific issues and effectuate corresponding enhancements. The pivotal role of these modules in ameliorating the IR pipeline and elevating system performance cannot be overstated. In this survey, we focus on the following four modules, which have been greatly enhanced by LLMs.
Query Rewriter is an essential IR module that seeks to improve the precision and expressiveness of user queries. Positioned at the early stage of the IR pipeline, this module assumes the crucial role of refining or modifying the initial query to align more accurately with the user’s information requirements. As an integral part of query rewriting, query expansion techniques, with pseudo relevance feedback being a prominent example, represent the mainstream approach to achieving query expression refinement. In addition to its utility in improving search effectiveness across general scenarios, the query rewriter finds application in diverse specialized retrieval contexts, such as personalized search and conversational search, thus further demonstrating its significance.
@@
Reader has evolved as a crucial module with the rapid development of LLM technologies. Its ability to comprehend real-time user intent and generate dynamic responses based on the retrieved text has revolutionized the presentation of IR results. In comparison to presenting a list of candidate documents, the reader module organizes answer texts more intuitively, simulating the natural way humans access information. To enhance the credibility of generated responses, the integration of references into generated responses has been an effective technique of the reader module.
Furthermore, researchers explore unifying the above modules to develop a novel LLM-driven search model known as the Search Agent. The search agent is distinguished by its simulation of an automated search and result understanding process, which furnishes users with accurate and readily comprehensible answers. WebGPT [24] serves as a pioneering work in this category, which models the search process as a sequence of actions of an LLM-based agent within a search engine environment, autonomously accomplishing the whole search pipeline. By integrating the existing search stack, search agents have the potential to become a new paradigm in future IR.
@@ 2.2 Large Language Models
The evolution of LMs can be categorized into four primary stages, as discussed in prior literature [52]. Initially, LMs were rooted in statistical learning techniques and were termed statistical language models. These models tackled the issue of word prediction by employing the Markov assumption to predict the subsequent word based on preceding words. Thereafter, neural networks, particularly recurrent neural networks (RNNs), were introduced to calculate the likelihood of text sequences and establish neural language models. These advancements made it feasible to utilize LMs for representation learning beyond mere word sequence modeling. ELMo [58] first proposed to learn contextualized word representations through pre-training a bidirectional LSTM (biLSTM) network on large-scale corpora, followed by fine-tuning on specific downstream tasks. Similarly, BERT [59] proposed to pre-train a Transformer [60] encoder with a specially designed Masked Language Modeling (MLM) task and Next Sentence Prediction (NSP) task on large corpora. These studies initiated a new era of pre-trained language models (PLMs), with the “pre-training then fine-tuning” paradigm emerging as the prevailing learning approach. Along this line, numerous generative PLMs (e.g., GPT-2 [33], BART [61], and T5 [57]) have been developed for text generation problems including summarization, machine translation, and dialogue generation. Recently, researchers have observed that increasing the scale of PLMs (e.g., model size or data amount) can consistently improve their performance on downstream tasks (a phenomenon commonly referred to as the scaling law [62, 63]). Moreover, large-sized PLMs exhibit promising abilities (termed emergent abilities [42]) in addressing complex tasks, which are not evident in their smaller counterparts. Therefore, the research community refers to these large-sized PLMs as large language models (LLMs).
@@ 3 Query Rewriter
Query rewriting in modern IR systems is essential for improving search query effectiveness and accuracy. It reformulates users’ original queries to better match search results, alleviating issues like vague queries or vocabulary mismatches between the query and target documents. This task goes beyond mere synonym replacement, requiring an understanding of user intent and query context, particularly in complex searches like conversational queries. Effective query rewriting enhances search engine performance.
Traditional methods for query rewriting improve retrieval performance by expanding the initial query with information from highly-ranked relevant documents. Mainly-used methods include relevance feedback [87, 88, 89, 90, 91, 92], word-embedding based methods [93, 94] etc. However, the limited ability of semantic understanding and comprehension of user search intent limits their performance in capturing the full scope of user intent.
Recent advancements in LLMs present promising opportunities to boost query rewriting capabilities. On one hand, given the context and subtleties of a query, LLMs can provide more accurate and contextually relevant rewrites. On the other hand, LLMs can leverage their extensive knowledge to generate synonyms and related concepts, enhancing queries to cover a broader range of relevant documents, thereby effectively addressing the vocabulary mismatch problem. In the following sections, we will introduce the recent works that employ LLMs in query rewriting.
Figure 4: An example of LLM-based query rewriting for conversational search. The example is cited from LLMCS [95]. The LLM is used to generate a query based on the demonstrations and previous search context. Additional responses are required to be generated for improving the query understanding. $N=0$ and $N>0$ correspond to zero-shot and few-shot scenarios.
3.1 Rewriting Scenario
Query rewriting typically serves two scenarios: ad-hoc retrieval, which mainly addresses vocabulary mismatches between queries and candidate documents, and conversational search, which refines queries based on evolving conversations. The upcoming section will delve into the role of query rewriting in these two domains and explore how LLMs enhance this process.
@@ 3.1.1 Ad-hoc Retrieval
In ad-hoc retrieval, queries are often short and ambiguous. In such scenarios, the main objectives of query rewriting include adding synonyms or related terms to address vocabulary mismatches and clarifying ambiguous queries to more accurately align with user intent. From this perspective, LLMs have inherent advantages in query rewriting. Primarily, LLMs have a deep understanding of language semantics, allowing them to capture the meaning of queries more effectively. Besides, LLMs can leverage their extensive training on diverse datasets to generate contextually relevant synonyms and expand queries, ensuring broader and more precise search result coverage. Additionally, studies have shown that LLMs’ integration of external factual corpora  [96, 97, 98, 99] and thoughtful model design [100] further enhance their accuracy in generating effective query rewrites, especially for specific tasks.
Currently, there are many studies leveraging LLMs to rewrite queries in adhoc retrieval. We introduce the typical method Query2Doc [86] as an example. As shown in Figure 3, Query2Doc prompts the LLMs to generate a relevant passage according to the original query (“when was pokemon green released?”). Subsequently, the original query is expanded by incorporating the generated passage. The retriever module uses this new query to retrieve a list of relevant documents. Notably, the generated passage contains additional detailed information, such as “Pokemon Green was released in Japan on February 27th”, which effectively mitigates the “vocabulary mismatch” issue to some extent.
In addition to addressing the “vocabulary mismatch” problem [101, 102, 96, 97, 98, 99], other works utilize LLMs for different challenges in ad-hoc retrieval. For instance, PromptCase [103] leverages LLMs in legal case retrieval to simplify complex queries into more searchable forms. This involves using LLMs to identify legal facts and issues, followed by a prompt-based encoding scheme for effective language model encoding.
@@ 3.1.2 Conversational Search
Query rewrites in conversational search play a pivotal role in enhancing the search experience. Unlike traditional queries in ad-hoc retrieval, conversational search involves a dialogue-like interaction, where the context and user intent evolve with each interaction. In conversational search, query rewriting involves understanding the entire conversation’s context, clarifying any ambiguities, and personalizing responses based on user history. The process includes dynamic query expansion and refinement based on dialogue information. This makes conversational query rewriting a sophisticated task that goes beyond traditional search, focusing on natural language understanding and user-centric interaction.
@@
The LLMCS framework [95] is a pioneering approach that employs LLMs to effectively extract and understand user search intent within conversational contexts. As illustrated in their work, LLMCS uses LLMs to produce both query rewrites and extensive hypothetical system responses from various perspectives. These outputs are combined into a comprehensive representation that effectively captures the user’s full search intent. The experimental results show that including detailed hypothetical responses with concise query rewrites markedly improves search performance by adding more plausible search intent. Ye et al. [104] claims that human query rewrite may lack sufficient information for optimal retrieval performance. It defines four essential properties for well-formed LLM-generated query rewrites. Results show that their method informative query rewrites can yield substantially improved retrieval performance compared to human rewrites.
Besides, LLMs can be used as a data expansion tool in conversational dense retrieval. Attributed to the high cost of producing hand-written dialogues, data scarcity presents a significant challenge in the domain of conversational search. To address this problem, CONVERSER [105] employs LLMs to generate synthetic passage-dialogue pairs through few-shot demonstrations. Furthermore, it efficiently trains a dense retriever using a minimal dataset of six in-domain dialogues, thus mitigating the issue of data sparsity.
3.2 Rewriting Knowledge
Query rewriting typically necessitates additional corpora for refining initial queries. Considering that LLMs incorporate world knowledge in their parameters, they are naturally capable of rewriting queries. We refer to these methods, which rely exclusively on the intrinsic knowledge of LLMs, as LLM-only methods. While LLMs encompass a broad spectrum of knowledge, they may be inadequate in specialized areas. Furthermore, LLMs can introduce concept drift, leading to noisy relevance signals. To address this issue, some methods incorporate domain-specific corpora to provide more detailed and relevant information in query rewriting. We refer to methods enhanced by domain-specific corpora to boost LLM performance as corpus-enhanced LLM-based methods. In this section, we will introduce these two methods in detail.
3.2.1 LLM-only methods